if there are 52 grams of sulphuric acid in a 500 gram solution, what is the percent strength of acid?
Answer by stanbon(75887)

if there are 52 grams of sulphuric acid in a 500 gram solution, what is the percent strength of acid?
-------------------------------
Ans: 52/500 = 104/1000 = 0.104 = 10.4%
---------------
Cheers,
Stan H.
------------